Reese, a calendar-year taxpayer, uses the cash method of accounting for her sole proprietorship. In late December, she received a $63,000 bill from her accountant for consulting services related to her small business. Reese can pay the $63,000 bill anytime before January 30 of next year without penalty. Assume Reese's marginal tax rate is 32 percent this year and 35 percent next year, and that she can earn an after-tax rate of return of 11 percent on her investments.

a. What is the after-tax cost if she pays the $63,000 bill in December?

b. What is the after-tax cost if she pays the $63,000 bill in January? Use Exhibit 3.1. (Do not round intermediate calculations. Round your answer to the nearest whole dollar amount.)

c. Should Reese pay the $63,000 bill in December or January?

d. What is the after-tax cost if she expects her marginal tax rate to be 24 percent next year and pays the $63,000 bill in January? Use Exhibit 3.1. (Round your answer to the nearest whole dollar amount.)

e. Should Reese pay the $63,000 bill in December or January if she expects her marginal tax rate to be 32 percent this year and 24 percent next year?
Answer #1
a. After-tax cost if Reese pays the $63,000 bill in December
=63000-32%(63000)
$42,840
b. After-tax cost if Reese pays the $63,000 bill in January
Consulting Cost $63,000.00
-Interest Income (1 month) -$577.50 =63000*11%/12
Net Cost $62,422.50
-Tax @ 35% -$21,847.88
After Tax Cost $40,575
c. Reese should pay bill in January as Net after tax cost is lesser in January.
d.
Consulting Cost $63,000.00
-Interest Income (1 month) -$577.50 =63000*11%/12
Net Cost $62,422.50
-Tax @ 24% -$14,981.40
After Tax Cost $47,441
e. Reese should pay bill in December as Net after tax cost is lesser in January.